OpenStudy (anonymous):

I need help with 4 questions please and thank you! Also giving medal!!! -------------------------------------------------------- 1. Select the point that is on the right side of the vertex of y = |x + 3| + 4. A) (–5, 6) B) (–4, 5) C) (–3, 4) D) (–1, 6) 2. What is the vertex of y = 2|x| – 1 ? A) (–1, 0) B) (0, –1) C) (0, 1) D) (1, 0) 3. What is the solution set for |x – 2|– 1 > 2? A) x > 5 or x < –1 B) x < 5 or x > –1 C) x > –5 or x < 1 D) x < 5 or x > –5

OpenStudy (anonymous):

4. Select the inequality of the graph shown below. A) y greater than or equal to |x + 2| + 1 B) y less than or equal to |x + 2| + 1 C) y greater than or equal to |x – 2| + 1 D) y less than or equal to |x – 2| + 1

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

how bout we do this one y = |x + 3| + 4 first

OpenStudy (anonymous):

Okay. Thanks.

OpenStudy (anonymous):

for the first coordinate of the vertex, set \(x+3=0\) and solve in one step what do you get?

OpenStudy (anonymous):

x = -3

OpenStudy (anonymous):

ok good, and the second coordinate of the vertex is that \(+4\) hanging out at the end

OpenStudy (anonymous):

so the vertex is \((-3,4)\)

OpenStudy (anonymous):

Oh Okay. That is it? Wow, thanks.

OpenStudy (anonymous):

this one What is the vertex of y = 2|x| – 1 is similar only easier, since if you set \(x=0\) there are no steps in solving, first coordinate of the vertex is \(0\) and the second coordinate is that \(-1\) hanging out at the end so it is \((0,-1)\) and yes, it is that easy

OpenStudy (anonymous):

Wow, thanks.

OpenStudy (anonymous):

this one |x – 2|– 1 > 2 start by adding 1 on both sides of the inequality to get \[|x-2|>3\]

OpenStudy (anonymous):

Okay, I am following.

OpenStudy (anonymous):

then you have to solve TWO inequalities separately, but both are easy you get \[x-2>3\] as one inequality, and \(x-2<-3\) as the other each is solved in one step what do you get?

OpenStudy (anonymous):

x < - 5 and x > 1???

OpenStudy (anonymous):

no, i think your mistake was you subtracted 2 from both sides, but because you have \(x-2\) you want to ADD 2 to both sides to get \(x\) by itself try again but this time add 2 to both sides

OpenStudy (anonymous):

I am confused. :/

OpenStudy (anonymous):

Wait.

OpenStudy (anonymous):

I got it.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

x<1 x>-2?

OpenStudy (anonymous):

No, that doesn't seem right. :/

OpenStudy (anonymous):

that is because you switched the inequalities by mistake

OpenStudy (anonymous):

\(x-2>3\) add 2 to both sides, leave the inequality alone

OpenStudy (anonymous):

x > 5

OpenStudy (anonymous):

ok good that is the solution to one of them now you still have to solve \[x-2<-3\] and again add 2 to both sides

OpenStudy (anonymous):

x < -1

OpenStudy (anonymous):

x > 5 and x < -1

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

YAY

OpenStudy (anonymous):

THANKYOU! (:

OpenStudy (anonymous):

to be picky and technical it is \(x<-1\) OR \(x>5\) not "and" since no number fits both categories

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (anonymous):

yw still one more to go right?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

#4.

OpenStudy (anonymous):

ok so as you can see from the picture, the shaded region is above the two lines, so you know it must be \(y\geq something\)

OpenStudy (anonymous):

that means you can ignore answer B and D and only consider A and C

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

It is A.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

lets go slow here it takes me a minute to scroll up and see the answers

OpenStudy (anonymous):

Okay. Sorry.

OpenStudy (anonymous):

A is \(y \geq|x + 2| + 1\) now lets find the vertex of \(y=|x+2|+1\)

OpenStudy (anonymous):

first coordinate of the vertex, you set \(x+2=0\) and solve what do you get ?

OpenStudy (anonymous):

did i lose you?

OpenStudy (anonymous):

No. I am here.

OpenStudy (anonymous):

x = -2

OpenStudy (anonymous):

Sorry.

OpenStudy (anonymous):

right, so the vertex of\(y=|x+2|+1\) is \((-2,1)\) but that is NOT the vertex in the picture, is it?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

you can see from the picture that the vertex, that point on the bottom, is \((2,1)\) and so the graph is of \(y=|x-2|+1\)

OpenStudy (anonymous):

So it is C?

OpenStudy (anonymous):

that means you are looking at \(y\geq |x-2|+1\)

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

yeah, it is always C

OpenStudy (anonymous):

Great thank you so much!!!!

OpenStudy (anonymous):

You're a life saver ! (:

OpenStudy (anonymous):

yw you got more or is that it?

OpenStudy (anonymous):

That is it for now. (: Thank you so much! I really appreciate it.

OpenStudy (anonymous):

no problem hope it was more or less clear in case you have to take a test or something